### v3

# Stand-In: A Lightweight and Plug-and-Play Identity Control for Video Generation

Abstract
Generating high-fidelity human videos that match user-specified identities is important yet challenging in the field of generative AI. Existing methods often rely on an excessive number of training parameters and lack compatibility with other AIGC tools. In this paper, we propose Stand-In, a lightweight and plug-and-play framework for identity preservation in video generation. Specifically, we introduce a conditional image branch into the pre-trained video generation model. Identity control is achieved through restricted self-attentions with conditional position mapping. Thanks to these designs, which greatly preserve the pretrained prior of the video generation model, our approach is able to outperform other full-parameter training methods in video quality and identity preservation, even with just $\sim$1% additional parameters and only 2000 training pairs. Moreover, our framework can be seamlessly integrated for other tasks, such as subject-driven video generation, pose-referenced video generation, stylization, and face swapping.
Figure 1: Given a reference image, our method generates videos with strong identity preservation. Furthermore, the framework’s plug-and-play design enables seamless integration into diverse applications for enhanced identity consistency.
footnotetext: Equal Contribution.footnotetext: # Corresponding Author.
1 Introduction
With the rapid advancement of diffusion models [NEURIPS2020_4c5bcfec, podell2023sdxl, Peebles_2023_ICCV], video generation [opensora, opensora2, kong2024hunyuanvideo, hong2022cogvideo] has become a pivotal aspect of generative AI. Among its diverse applications, identity-preserving video generation holds profound significance. The goal of this task is to generate high-quality videos that consistently maintain the identity of a given reference image containing a face. It has widespread utility across film, advertising and gaming industries, etc.
Existing methods can be roughly classified into two categories. Early methods [ID-Animator, consistid] use an explicit face encoder for identity feature extraction, while recent methods [hu2025hunyuancustom, liu2025phantom] fully train the diffusion transformer. However, face-encoder-based methods lack flexibility and struggle to capture fine facial details essential for high-quality video generation. The full-parameter training methods require huge training resources and lack compatibility with other applications. Achieving robust identity preservation in a lightweight way remains critical yet challenging.
To overcome these limitations, we leverage the pre-trained VAE from the video generation model itself, enabling the conditional image to be mapped directly into the same latent space as the video. This approach naturally utilizes the model’s inherent capabilities to extract rich and detailed facial features, offering a more integrated and effective solution. Specifically, we employ restricted self-attention with conditional position mapping to merge the features of the reference image into the video. On the one hand, by preserving the core functionality of self-attention and the pretrained prior of the video generation model, our method achieves the highest facial similarity and naturalness in identity-preserving video generation with only $\sim$1% additional parameters and 2000 training pairs, which is shown in Figure 2. On the other hand, our method does not alter the architecture of the main video generation model and thus can be used in a plug-and-play manner for other applications. As shown in Figure 1, our framework can be extended to various tasks, including subject-driven generation, video stylization, and face swapping, all guaranteeing identity consistency. Additionally, benefit from integrating compatibility with VACE [vace], our approach significantly enhances facial similarity in pose-guided video generation. Our main contributions can be summarized as follows:
  • We present Stand-In, a lightweight and plug-and-play framework designed for identity-preserving video generation. By incorporating and training just $\sim$1% additional parameters, our approach achieves SOTA results in identity preservation, video quality, and prompt following.
  • To inject identity information without explicit face feature extractors, we introduce a conditional image branch to the video generation model. The image and video branches share information through restricted self-attention with conditional position mapping. With these designs, identity preservation can be learned well with a small dataset.
  • The proposed framework exhibits high compatibility and generalizability. Although trained only on real-people data, our method generalizes to other subjects, such as cartoons and objects. Moreover, our method can be plug-and-play applied to other tasks, such as pose-guided video generation, video stylization and face swapping.
Figure 2: Comparison with SOTA identity-preserving video generation methods. The size of bubbles represents the number of need-to-train parameters for identity preservation. Our approach achieves the highest performance in both face similarity and naturalness, while utilizing the fewest parameters.
2 Related Work
Video Generation
Current video generation models are predominantly built on diffusion frameworks [NEURIPS2020_4c5bcfec], with an evolution in architecture from U-Net-based designs [blattmann2023stablevideodiffusionscaling] to DiT-based approaches [kong2024hunyuanvideo, wan2.1, ma2025latte, DiT, wang2023laviehighqualityvideogeneration, VideoCrafter2, T2V-Turbo]. In the era of U-Net-based diffusion models, text-to-image (T2I) frameworks [rombach2021highresolution, podell2023sdxl] were extended to video generation by introducing 3D convolutions and temporal attention  [blattmann2023stablevideodiffusionscaling]. AnimateDiff [guo2023animatediff] further advanced this direction by reusing pre-trained text-to-image model weights to leverage their strong spatial generation capabilities by adding temporal layers. Latte [ma2025latte] introduced a spatial-temporal separation mechanism, assigning distinct DiT blocks to process spatial and temporal information independently. This approach was later replaced by 3D full attention mechanisms, which offered more integrated processing. CogVideoX [yang2024cogvideox] and HunyuanVideo [kong2024hunyuanvideo] combined 3D-VAE  [yu2023language] with MM-DiT [esser2024scaling] to enhance video generation capabilities. WAN2.1 [wan2.1] employs a 3D-VAE and adopts a DiT backbone for denoising, injecting semantic information into the diffusion process through cross-attention.
Identity-Preserving Generation
Prior to the advent of zero-shot identity-preserving algorithms, generating content with consistent identity typically relied on case-by-case fine-tuning.[Dreambooth, Lora, textual_inversion, multi-concept, dreamvideo, customvideo, motionbooth], In contrast, training-free identity-preserving image generation methods[ye2023ip, shi2024instantbooth, guo2025pulid, wang2024instantid, he2024uniportrait, sara2025ipcompose, qian2024omni, han2024face, mou2025dreamounifiedframeworkimage] offer a zero-shot personalization solution by integrating identity features into pre-trained foundation models. These methods typically introduce parameterized plug-in modules or adapters to adjust and inject identity features into the generation process. A popular solution is to use a Face Encoder to extract face embeddings and inject them into the generation process via cross-attention. For example, InstantID[wang2024instantid] and PuLID[guo2025pulid] can generate high-quality, identity-preserving images.
In the field of identity-preserving video generation. Early methods commonly rely on explicit face encoders for facial feature extraction to generate videos with identity-preserving. ID-animator [ID-Animator] leverages a pre-trained text-to-video diffusion model in conjunction with a lightweight face adapter to encode ID-relevant embeddings from adaptable facial latent queries. ConsistID [consistid] aims to maintain identity consistency through frequency decomposition in diffusion transformer. Phantom [liu2025phantom] can also preserve identity consistency in the human domain as a unified subject-consistent video generation framework. HunyuanCustom [hu2025hunyuancustom] is a multi-modal customized video generation framework that emphasizes identity consistency while supporting diverse input modalities. By introducing advanced condition injection mechanisms and identity-preserving strategies, it achieves excellent performance in high-quality video generation. They employed a full fine-tuning for the diffusion transformer, requiring huge training resources.
3 Method
In this section, we first introduce the overall framework of the proposed method in Section 3.1. Next, we detail the restricted self-attention mechanism in Section 3.2 and conditional position mapping in Section 3.3. Finally, we present the data collection process in Section 3.4.
3.1 Overall Framework
To extract facial features, early methods [ID-Animator, consistid] rely on explicit face encoders, which lack flexibility and often fail to preserve fine facial details critical for high-quality reconstruction. In contrast, we propose using the pre-trained VAE from the video generation model. This strategy maps the conditional image directly into the same latent space as the video and allows us to naturally take advantage of the built-in ability of the pre-trained video generation model to extract rich facial features.
The overall framework is illustrated in Figure 3. We use Wan2.1-14B-T2V [wan2.1] as the video generation base model, which adopts a Diffusion Transformer (DiT) architecture. Given a reference image containing a face, we first encode it into the latent space using the pre-trained VAE encoder. The image latents undergo the same patchification and encoding procedures as the video latents. Then, the image tokens are concatenated with video tokens along the sequence dimension and processed jointly through successive blocks. Finally, image tokens are discarded at the final layer.
[…]
3.2 Restricted Self-Attention
Figure 3: The overview of our identity-preserving text-to-video generation framework. We introduce a conditional image branch alongside the original video branch. Given the conditional image, the VAE encoder maps it into tokens, which are concatenated with the video latent tokens and then sent to the DiT. Within the DiT blocks, identity information is incorporated into the video features through restricted self-attention.
Figure 4: Design of our Restricted Self-Attention: For the input video and image tokens, we compute their Query, Key, and Value matrices independently. Next, we apply Conditional Position Mapping to the Query and Key matrices. Finally, the image matrices operate independently, while the video Query performs attention using the concatenation of the image and video Key and Value matrices.
In the aforementioned DiT blocks, reference image and video tokens are processed independently through most modules (including layer normalization, cross-attention, and feed-forward networks), except for the self-attention layer. The self-attention layer enables information exchange among all tokens, naturally allowing video tokens to refer to identity information. One intuitive solution is to direct concatenate the reference-image tokens with the generated video tokens, and pass them through the Vanilla Self-Attention. However, this approach has two main drawbacks. First, since the reference image serves as a static condition, it should remain unaffected by the dynamic contents of the video. Vanilla Self-Attention lets image queries attend to video contents, making it challenging to maintain the identity. Second, this joint self-attention provides no guarantee that video tokens will actually refer to the image tokens. The model may ignore the reference image and generate scenes without the target identity, which is shown in Figure 5.
Therefore, to incorporate identity information while preserving its independence, we propose replacing the vanilla self-attention layers in DiT with a restricted version that explicitly prevents image queries from attending to video keys. As shown in Figure 4, for a self-attention layer, we first independently compute the Query, Key, and Value for image and video tokens, denoted as $Q_{I}$, $K_{I}$, $V_{I}$ and $Q_{V}$, $K_{V}$, $V_{V}$ respectively. Then, we concatenate $K_{V}$ with $K_{I}$ and $V_{V}$ with $V_{I}$ for $Q_{V}$. To enhance the model’s ability to utilize identity-related information while preserving its inherent generative robustness, we incorporate Low-Rank Adaptation (LoRA) into the QKV projection of image tokens. For analysis, we visualize attention maps specifically for video queries attending to the reference image by averaging attention maps across all DiT blocks. As shown in Figure 4, in contrast to the vanilla baseline, our Restricted Self-Attention concentrates attention on facial regions of the reference image and yields prompt-faithful frames that preserve the subject’s identity.
Given that the timestep for the conditional image is fixed at $s_{ref}=0$, its Key and Value matrices remain constant throughout the diffusion denoising process. Therefore, during inference, we can cache $K_{I}$ and $V_{I}$ to accelerate computation, named KV Caching. These matrices are computed and stored during the first denoising step, eliminating the need for redundant recalculations in subsequent steps.
3.3 Conditional Position Mapping
To effectively differentiate image and video tokens in the restricted self-attention, we use a specialized conditional position mapping strategy. Specifically, we employ 3D Rotary Positional Embedding (RoPE) [RoPE], where all tokens associated with the reference image are assigned a distinct and dedicated coordinate space. This ensures clear separation and facilitates precise modeling of interactions between the reference image and video tokens.
For the temporal dimension, we assign a fixed temporal index of -1 to the reference image tokens, while mapping video tokens to nonnegative temporal positions. This assignment establishes image tokens as temporally invariant conditional inputs. In this way, the model is encouraged to treat the identity information from the reference image as a constant guide throughout the entire denoising process, rather than conflating it with transient, frame-specific features in the temporal sequence.
For spatial dimensions, we employ a disjoint coordinate strategy to enforce spatial decoupling between reference image and video content. While video frames occupy coordinates within the domain $(h,w)\in[0,H_{V})\times[0,W_{V})$, we map reference image tokens to a dedicated coordinate subspace $H_{V},H_{V}+H_{I})\times[W_{V},W_{V}+W_{I})$, where $H_{I}$ and $W_{I}$ represent the reference image dimensions.
Figure 5: Effect of Restricted Self-Attention (RSA). Given the reference image and the prompt (left column), we visualize attention map to reference-image tokens. Under Vanilla Self-Attention (top row), attention diffuses into background regions and the output skews toward a garden scene. With RSA (bottom row), attention concentrates on facial regions, maintaining the subject’s identity.
Denoting $p_{I}$ as the coordinate for image tokens and $p_{V}$ for video tokens, we apply 3D RoPE as follows:
$Q^{\prime}_{I}$ $=Q_{I}\cdot p_{I},$ $K^{\prime}_{I}$ $=K_{I}\cdot p_{I},$ (1)
$Q^{\prime}_{V}$ $=Q_{V}\cdot p_{V},$ $K^{\prime}_{V}$ $=K_{V}\cdot p_{V}.$ (2)
where $\cdot$ denotes the Hadamard product. The restricted self-attention outputs are computed as:
$\text{Out}_{I}$ $=\text{Attention}(Q^{\prime}_{I},K^{\prime}_{I},V_{I}),$ (3)
$\text{Out}_{V}$ $=\text{Attention}(Q^{\prime}_{V},[K^{\prime}_{V},K^{\prime}_{I}],[V_{V},V_{I}]),$ (4)
where $\cdot$$\cdot$] denotes concatenation.
Figure 6: Effect of Conditional Position Mapping (CPM). Compared with Shared Postion Mapping, our CPM, where the reference tokens are mapped to a disjoint spatial space, better preserves the pretrained positional prior and yields more stable scenes.
This non-overlapping spatial allocation achieves two main goals through geometric separation. By separating the reference tokens from the video coordinate grid, the design naturally reduces false spatial correlations and better preserves the backbone’s pretrained positional prior. Compared with Shared Position Mapping, our Conditional Position Mapping generates more reliable video with more stable scene layouts, which is shown in Figure 6. At the same time, this separate coordinate system maintains the reference image’s semantic meaning by keeping it as a global identity prior. Consequently, the model is guided to focus on extracting overall semantic features from the reference tokens, rather than treating them as spatially localized patterns that need to align positionally with the video content.
3.4 Dataset Collection and Processing
We construct a human-centric video dataset containing 2,000 high-resolution sequences from publicly available sources. The dataset guarantees a diverse and comprehensive representation, comprising various ethnic groups, age ranges, gender identities, and a wide array of actions. Using the VILA [lin2024vila] multimodal captioning framework, we automatically generate dense textual annotations for each video, establishing strong text-video alignment.
[…]
For each video clip, the corresponding reference facial image is extracted from the original, pre-resampled video. The extraction pipeline is as follows:
  1. 1.
    5 frames are randomly selected from the original video.
  2. 2.
    Faces are detected and cropped using RetinaFace[Deng2020CVPR].
  3. 3.
    The cropped face images are resized to $\times512$ pixels.
  4. 4.
    BiSeNet [BiseNet] is used for face parsing, and the background is replaced with a solid white color to prevent any leakage of background information.
Figure 7: Examples from our human-centric video dataset.
Examples of the final image-text-video pairs for our training can be found in Figure 7.
Table 1: Quantitative comparison with state-of-the-art identity-preserving video generation methods. We evaluate across three key metrics: Face Similarity, Naturalness, and Prompt Following. For all metrics, higher values indicate better performance. The best and second-best results in each column are highlighted in bold and underlined, respectively.
Method
Face-
Similarity
$\uparrow$
Naturalness $\uparrow$
Prompt-
Following
$\uparrow$
Closed-Source Methods
Kling 0.410 3.900 19.921
Hailuo 0.577 3.750 20.649
Pika-2.1 0.323 3.644 20.649
Vidu-2.0 0.361 3.600 18.998
Open-Source Methods
ID-Animator [ID-Animator] 0.316 3.211 16.677
SkyReels-A2-P14B [SkyReels] 0.546 3.411 19.110
EchoVideo [EchoVideo] 0.487 3.456 19.263
ConcatID-CogVideoX [Concat-ID] 0.439 3.372 19.359
ConcatID-WAN [Concat-ID] 0.501 3.650 19.671
Hunyuan-Custom [HunyuanCustom] 0.622 3.367 19.853
ConsistID [consistid] 0.432 3.233 20.552
VACE-P1.3B [vace] 0.180 3.567 20.591
VACE-1.3B [vace] 0.223 3.611 20.527
VACE-14B [vace] 0.647 3.728 19.520
Phantom-1.3B [liu2025phantom] 0.440 3.567 20.364
Phantom-14B [liu2025phantom] 0.519 3.828 20.476
Stand-In (Ours) 0.724 3.922 20.594
Figure 8: Comparison on identity-preserving video generation. Please refer to the supplementary material for full prompts.
4 Experiments
4.1 Implementation Details
We adopt LoRA with rank 128, applied only to the QKV projections for image tokens in each DiT block. For the 14B-parameter Wan2.1 model, this adds just 153M trainable parameters (1% of the base model), increasing feed-forward time by 3.6% and FLOPs by 2.6%. During inference with KV caching, overhead is minimal: runtime rises by only 2.3% and FLOPs by 0.07% compared to the video generation base model. This negligible cost shows our identity-preserving method is lightweight.
Table 2: User study results for subjective evaluation. The best and second-best results in each column are highlighted in bold and underlined, respectively.
Method Face Similarity $\uparrow$ Video Quality $\uparrow$
Hunyuan-Custom[HunyuanCustom] 3.34 2.92
VACE-14B [vace] 3.00 3.07
Phantom-14B [liu2025phantom] 2.37 2.92
ConsistID [consistid] 2.25 2.46
Kling 2.21 3.09
Stand-In (Ours) 4.10 4.08
The model is trained over 3000 steps on Nvidia H20 GPUs with a batch size of 48. For inference, BiSeNet is adopted as an automatic preprocessing step.
Figure 9: Our model generalizes to unseen ordinary individuals across diverse ethnicities and age groups, despite being trained with only $\sim$1% additional parameters and just 2000 training pairs.
Figure 10: Our results on subjects other than real-person. Please refer to the supplementary material for full prompts.
Figure 11: Comparison on pose-guided video generation against VACE.
Figure 12: Application of our model in video face swapping.
Figure 13: Our model applied with stylization LoRA.
4.2 Quantitative Analysis
To evaluate identity preservation and visual quality, we use the two most important and heavily weighted evaluation metrics from the OpenS2V benchmark [yuan2025opens2vnexusdetailedbenchmarkmillionscale]: facial similarity and naturalness. To evaluate the relevance between generated video and textual description, we use X-CLIP [xclip], a pre-trained video-text multimodal model. Results are shown in Table 1 and Figure 8. As shown in Figure 9, despite being trained with only $\sim$1% additional parameters and just 2000 training pairs, our model also demonstrates strong generalization across individuals of diverse ethnicities and age groups.
Face Similarity
This metric evaluates the model’s ability to maintain identity consistency. It is calculated as the average cosine similarity between the CurricularFace embeddings of the reference image and the faces detected in the generated video frames. As shown in Table 1, our proposed method, Stand-In, achieves a score of 0.724, outperforming all other compared methods. This result demonstrates the effectiveness of our approach in generating facial features that remain highly consistent with the source identity.
Naturalness
This metric is primarily designed to evaluate the naturalness of the generated videos. It leverages GPT-4o to approximate human judgment of video realism, taking into account factors such as physical plausibility and the absence of noticeable AI artifacts. Following the OpenS2V protocol, a holistic score ranging from 1 to 5 is assigned. In this evaluation, our method achieves a score of 3.922, demonstrating that the improvements in identity fidelity are achieved without compromising the overall visual realism of the generated videos.
Prompt Following
As shown in Table 1, our method ranks second among all compared methods and first among open-source methods. This result demonstrates that our identity-preserving ability can be achieved without hurting the prompt-following performance.
4.3 User Study
The user study involves 20 participants. We randomly select 10 test videos from the benchmark and ask the participants to rate each video across two dimensions: facial similarity and video quality. The latter dimension encompasses aspects such as naturalness, visual aesthetics, and alignment with the provided text descriptions. Ratings are given on a 5-point scale (1 to 5), and the final scores for each dimension are obtained by averaging the ratings across all participants and test videos. As shown in Table 2, our method outperforms the comparative methods.
4.4 Plug-and-Play to Other Applications
Subject-Driven Video Generation
Although trained only with human data, our framework can be zero-shot applied to non-human subjects without any additional fine-tuning. This is because we use the pretrained VAE and video generation model to extract rich features, and learn alignment through paired data and an effective attention mechanism. This zero-shot ability can hardly be achieved by traditional identity-preserving methods relying on face encoders. As shown in Figure 10, our method exhibits strong subject consistency on the teddy bear object and the cartoon character.
Pose-Guided Video Generation
The proposed conditional image branch is designed based on the LoRA module, which ensures inherent compatibility with other DiT-based models. To validate this, we conducted experiments on the pose-guided video generation task using the VACE framework [vace]. As illustrated in Figure 11, integrating our method significantly improves the facial identity similarity in the generated videos. This not only demonstrates the plug-and-play nature of our approach but also highlights its robustness in preserving identity consistency.
Video Stylization
By applying our framework in conjunction with video stylization LoRA, we demonstrate its ability to achieve effective style transfer while maintaining strong identity consistency. As shown in Figure 13, our method successfully renders the artistic style as well as preserving the facial features of the reference image, further showing its versatility and robustness.
Video Face Swapping
Our framework is also capable of video face swapping, which can be achieved via zero-shot inpainting [inpainting]. Figure 12 shows that our method not only achieves high-quality facial identity transfer but also maintains strong temporal consistency across frames, resulting in coherent and high-quality videos.
4.5 Ablation Study
Effectiveness of Restricted Self-Attention (RSA)
To verify the role of RSA, we replaced it with Vanilla Self-Attention (VSA), which concatenates the reference-image and video tokens and allows full bidirectional information flow. Because VSA treats the reference tokens as part of the dynamic video context, the model often fails to refer to the reference image. As shown in Figure 5, this naive design leads to diffused attention and unreliable identity preservation. In contrast, RSA constrains attention so that image queries cannot attend to video tokens. This design keeps the reference representation static while still providing identity cues. As shown in Table 3, RSA improves face similarity from 0.422 to 0.724 and also slightly enhances naturalness, confirming its essential role in maintaining both identity consistency and visual quality.
Table 3: Ablation study on the core components of our method. Replacing Restricted Self-Attention (RSA) with Vanilla Self-Attention (VSA) or Conditional Position Mapping (CPM) with Shared Position Mapping (SPM) degrades both Face Similarity and Naturalness.
Method Face Similarity $\uparrow$ Naturalness $\uparrow$
RSA $\rightarrow$ VSA 0.422 3.855
CPM $\rightarrow$ SPM 0.536 3.755
Full Model 0.724 3.922
Effectiveness of Conditional Position Mapping (CPM)
We further investigate the effect of CPM by Replacing CPM with a Shared Position Mapping (SPM), which places reference and video tokens within the same positional coordinate system. As illustrated in Figure 6, SPM breaks the pretrained positional prior of the backbone model, resulting in unstable spatial layouts. By geometrically separating the coordinate spaces, CPM preserves the pretrained positional prior while maintaining the reference as a global identity prior. In Table 3, CPM increases face similarity from 0.536 to 0.724 and improves naturalness from 3.755 to 3.922, demonstrating its effectiveness in stabilizing video synthesis and enhancing identity fidelity.
5 Conclusion
We propose Stand-In, a lightweight, plug-and-play framework for high-fidelity, identity-preserving video generation. We introduce a conditional image branch into a pre-trained video generation model, and propose a restricted attention mechanism with conditional positional encoding to enable cross-branch information exchange. Despite training only 1% of the model’s additional parameters on a limited dataset of 2,000 pairs, our approach achieves high-quality video generation while maintaining strong identity fidelity. Experimental results demonstrate that Stand-In achieves state-of-the-art performance in identity-preserving text-to-video generation. Furthermore, it exhibits excellent performance on other tasks, including pose-guided video generation, stylization, and face swapping, proving its strong compatibility and broad application potential.
Stand-In: A Lightweight and Plug-and-Play Identity Control
for Video Generation
     Supplemental Material
6 Pseudocode
To more clearly and systematically present the core inference mechanism of the Stand-In method, we formalize its complete inference procedure in Algorithm 1. This pseudocode illustrates how the model jointly leverages the reference image during the diffusion process to preserve identity consistency, thereby enabling high-quality, identity-preserving video generation.
Algorithm 1 Inference Process of Stand-In
Input: Reference image tokens $T_{I}$, initial video latents $T_{V_{N}}$, list of timesteps $S=\{s_{N},s_{N-1},...,s_{1}\}$, DiT model $\mathcal{M}$.
Parameter: Zero timestep for the reference image, $s_{ref}=0$; Positional Encodings $p_{I},p_{V}$.
Output: Denoised video latents $T_{V_{0}}$.
1: Initialize an empty cache for each DiT block: $C_{k},C_{v}\leftarrow\{\},\{\}$
[…]
3: Let $is\_first\_step\leftarrow\text{true}$
4:for $s$ in $S$ do
5:  for each block $\mathcal{B}$ in model $\mathcal{M}$ do
6:   if is_first_stepis\_first\_step then
7:    Compute $(Q_{I},K_{I},V_{I})\leftarrow\mathcal{B}_{\text{before\_self-attn}}(T_{I},s_{ref})$
8:    Store $K_{I},V_{I}$ in cache: $C_{k}[\mathcal{B}]\leftarrow K_{I}$, $C_{v}[\mathcal{B}]\leftarrow V_{I}$
9:   else
10:    Retrieve from cache: $K_{I},V_{I}\leftarrow C_{k}[\mathcal{B}],C_{v}[\mathcal{B}]$
11:   end if
12:   Compute $(Q_{V},K_{V},V_{V})\leftarrow\mathcal{B}_{\text{before\_self-attn}}(T_{V},s)$
13:   // Self-attention start
14:   $(Q_{V}^{\prime},K_{V}^{\prime})\leftarrow\textnormal{{ApplyRoPE}}((Q_{V},K_{V}),p_{V})$
15:   $(Q_{I}^{\prime},K_{I}^{\prime})\leftarrow\textnormal{{ApplyRoPE}}((Q_{I},K_{I}),p_{I})$
16:   $K^{\prime}_{\text{cat}}\leftarrow\text{Concat}(K_{V}^{\prime},K_{I}^{\prime})$, $V_{\text{cat}}\leftarrow\text{Concat}(V_{V},V_{I})$
17:   $\text{Out}_{I}\leftarrow\text{Attention}(Q_{I}^{\prime},K_{I}^{\prime},V_{I})$
18:   $\text{Out}_{V}\leftarrow\text{Attention}(Q_{V}^{\prime},K^{\prime}_{\text{cat}},V_{\text{cat}})$
19:   // Self-attention end
20:   $(T_{V},T_{I})\leftarrow\mathcal{B}_{\text{after\_self-attn}}(\text{Out}_{V},\text{Out}_{I})$
21:  end for
22:  $is\_first\_step\leftarrow\text{false}$
23:  Predict denoised $T_{V}$ using scheduler and update for next step.
24:end for
25:return $T_{V_{0}}$
7 More Comparison Results
We provide more comparisons of our model with other identity-preserving video generation methods in Figures 14-17.
Figure 14: Visual comparison with other cutting-edge methods. Our method, Stand-In, demonstrates superior identity fidelity and detail preservation (1/4).
Figure 15: Visual comparison with other cutting-edge methods. Our method, Stand-In, demonstrates superior identity fidelity and detail preservation (2/4).
Figure 16: Visual comparison with other cutting-edge methods. Our method, Stand-In, demonstrates superior identity fidelity and detail preservation (3/4).
Figure 17: Visual comparison with other cutting-edge methods. Our method, Stand-In, demonstrates superior identity fidelity and detail preservation (4/4).
8 More Ablation Study Results
To further validate the effectiveness of our proposed Restricted Self-Attention (RSA) and Conditional Position Mapping (CPM), we provide additional qualitative comparisons in Figure 18 . The ablation focuses on two alternative configurations: (1) replacing RSA with vanilla self-attention (VSA), and (2) replacing CPM with Shared Position Mapping (SPM).
When RSA is replaced with Vanilla Self-Attention (VSA), the model suffers from severe identity drift. This confirms that unconstrained bidirectional attention causes the reference-image tokens to be contaminated by dynamic video content, preventing stable identity extraction. Similarly, when CPM is replaced with Shared Position Mapping (CPM → SPM), sharing the positional space results in clearly distorted facial proportions (e.g., enlarged face in the second example), revealing that the pretrained positional prior is severely disrupted. In contrast, our full model consistently maintains both accurate identity and natural facial geometry for both cases, demonstrating that RSA and CPM are jointly crucial for stable, identity-preserving video generation.
Figure 18: Visual results of the ablation study.
9 More Generalization Results
To further evaluate the robustness and generalization ability of our approach, Figure 19 showcases identity-preserving video generation results for a diverse set of subjects. Despite significant variations in age, ethnicity, skin tone, and facial geometry, our method consistently reconstructs key identity traits and maintains stable facial structure across frames. The generated videos exhibit natural motion, coherent expressions, and high perceptual fidelity. The text prompts from top to bottom are:
(1) A white man stands in front of his desk, facing the camera while speaking and gesturing to emphasize his points. He then picks up a pen and writes a few lines on a sheet of paper before looking back up with a confident smile, as if showing his progress to the viewer. The lighting is soft, giving him a natural and self-assured presence.
(2) An elderly Asian woman sits on a wooden balcony chair, smiling gently at the camera. She takes a small sip from her teacup, then looks up with a warm smile and nods. She reaches to adjust the shawl draped over her shoulders and closes her eyes slightly to enjoy the sunlight. Her expression is peaceful and kind, giving a sense of warmth and comfort.
(3) An elderly Black man sits on his home balcony, looking directly at the camera while holding a warm cup of tea with both hands. A gentle breeze brushes past as he softly blows across the tea’s surface and looks up with a warm smile. He then raises his cup and nods toward the camera in a calm and friendly gesture.
(4) A white woman faces the camera, gently sweeping a makeup brush across her face. She then offers a radiant smile, tilting her head slightly to show her makeup before returning to the front and giving a playful wink. Her movements are graceful and natural, complemented by soft lighting.
(5) An Asian man faces the camera, lightly adjusting the collar of his shirt before running his hand through his hair with a smile. He then raises a cup of tea and makes a small “cheers” gesture toward the camera. His expression is natural and confident, and his movements feel relaxed and smooth.
(6) A young Black woman smiles at the camera while adjusting her curly hair. She gently shakes a coffee cup toward the lens, as if sharing her morning energy with the viewer. She blinks playfully, makes an “OK” gesture, and smiles as sunlight from the window falls softly across her face, giving her a confident and natural expression.
(7) A young white boy smiles at the camera, giving a small wave. He glances down with a shy laugh before looking up again with a bright and innocent expression. Sunlight illuminates his face, making him look natural and adorable.
(8) A young Asian girl faces the camera and makes a heart shape with her hands. She then picks up a glass of juice, takes a sip, and smiles as she wipes the corner of her mouth. A breeze lifts her hair as she playfully brushes it aside, her eyes bright and expressive.
(9) A young Black girl smiles brightly at the camera. She waves hello softly, then forms a heart shape with her hands. Her eyes are bright, her smile sweet, and the sunlight behind her creates a warm and gentle atmosphere.
Figure 19: Given a single reference image (left), our method generates videos (right) that maintain strong identity consistency across diverse ethnicities and age groups, despite being trained with only $\sim$1% additional parameters and just 2000 training pairs.
10 Complete Video Generation Prompts
In this section, we present the complete prompts for Figures in the main paper.
Figure 8: Comparison on identity-preserving video generation
Left Panel:
The video features a man standing at an easel, focused intently as his brush dances across the canvas. His expression is one of deep concentration, with a hint of satisfaction as each brushstroke adds color and form. He wears a paint-splattered apron, and his hands move with confident precision. The setting, filled with scattered art supplies, open paint tubes, and unfinished sketches pinned to the wall, suggests an artist’s studio. A large window on one side allows sunlight to stream in, casting a soft glow across the room and illuminating the colors on his canvas. The atmosphere is creative and inspired, with the man’s intense focus and the lively colors on the canvas indicating a moment of artistic passion and expression.
Right Panel:
The video features a man with dark-haired hair, wearing a blue tank top and holding a pink tank top on a hanger. he appears to be in a clothing store or a similar retail environment, as there are racks of clothes visible in the background. The man is speaking to the camera, possibly providing a review or discussing the tank top he is holding. He has colorful bracelets on his wrist and is wearing a necklace with multiple beads. his expression suggests he is engaged in a conversation or presentation. The setting seems to be indoors, with artificial lighting illuminating the scene.
Figure 10: Our results on subjects other than real-person
Left Panel:
Shot in a medium shot of a brightly lit room, a girl, approximately seven or eight years old, stands in the center. She has long black hair and wears a light blue dress, her expression focused and gentle. Holding a doll in both hands, she presents her beloved toy to the camera. As the camera slowly zooms in, the details of her face are clearly visible: the soft fabric, the delicate stitching, and the slightly upturned corners of her mouth are all captured. The entire scene is filled with childlike innocence and warmth.
Right Panel:
The video features an anime girl standing on a busy street, surrounded by a hurried crowd. The buildings and shops in the background create a classic cityscape. The girl smiles as she puts her headphones on, her movements smooth and natural. Her expression is playful and relaxed, as if she’s about to immerse herself in her favorite music. The camera focuses on her face, capturing her joyful expression and vibrant energy. The background is slightly blurred, emphasizing the contrast between her and her surroundings, creating a sense of relaxed urban living.
Figure 12: Application of our model in video face swapping
Top Panel:
A woman is seated comfortably at a desk, facing the camera directly as if engaged in a video call with a friend or family member. Her gaze is intent and warm, complemented by a natural smile. The background consists of a personalized space, with photographs and a world map adorning the wall, thereby communicating a feeling of intimate and contemporary connection.
Bottom Panel:
A man is seated in a comfortable armchair, holding a thick, aged book on his lap. Drawn by the camera’s presence, he looks up, his face bearing a gentle and curious smile as if mildly interrupted. Soft, lateral lighting illuminates the scene. He is surrounded by floor-to-ceiling bookshelves, creating a tranquil atmosphere enveloped by knowledge and narrative.
11 More Results
Figure 20: We present more visual results of Stand-In (1/2).
Figure 21: We present more visual results of Stand-In (2/2).